# A UD Treebank for Bohairic Coptic

### v1

@@ 1 Introduction
Expanding digital corpora to include Bohairic texts presents substantial challenges, as existing tools for even basic preprocessing operations, such as word segmentation and lemmatization, let alone syntactic parsing, require tools trained on dialect-specific data (see Section 5 for some evaluation). Although Sahidic and Bohairic are dialectal manifestations of a single language—sharing a broadly consistent grammatical architecture and much of their lexical inventory, their phonological, orthographic and morphosyntactic systems diverge in ways that make NLP tools trained on Sahidic unsuitable for processing Bohairic texts.
At the most elementary level, the two dialects diverge already in their orthographic systems. Unlike earlier Egyptian – written in hieroglyphic and hieratic scripts over the preceding three millennia – Coptic adopted a modified version of the Greek alphabet, with additional characters of Demotic origin (ultimately derived from hieroglyphs) to represent sounds absent from Greek. Both dialects make use of six such additional letters, including the letter hore ({coptic}h1, Unicode U+03E9 lowercase / U+03E8 uppercase) for the voiceless guttural fricative /h/, but Bohairic distinguishes hore /h/ from khei /x/ ({coptic}j , Unicode U+2CC9 lowercase / U+2CC8 uppercase). This distinction is semantically consequential - while the Sahidic word {coptic}eh1rai (ehrai) can confusingly mean both ‘up’ or ‘down’, in Bohairic we see distinct forms that had merged in Sahidic: {coptic}ejrhi (exrēi) ‘down’ and {coptic}eh1rhi (ehrēi) ‘up’.
As an agglutinative language, Coptic combines multiple morphosyntactic elements into units known as bound groups. Following modern editorial conventions, these groups are defined by the presence of a single stressed lexical item at their core Layton (2011). Material that would normally be tokenized into separate words in annotated corpora often appears conflated into a single space-delimited string in Coptic. Such fusion is common, for example, in noun phrases or prepositional phrases, as in 1, or in auxiliaries and clitics attaching to verbs, as in 1, both examples in the Bohairic dialect.
@@
In this paper we use hyphens here to split such space-delimited word forms into tokens, which correspond to units that can receive independent parts of speech, such as nouns and verbs, articles and prepositions, etc., though such separators do not exist in source texts and are represented in the treebank directly through tokenization. For words containing derivational affixes as in 1, and compound nouns or verbs 1, we additionally provide a segmentation into component parts in an additional annotation called MSeg in the conllu format MISC field, following existing practices in UD treebanks. [^1] See https://universaldependencies.org/misc.html#mseg
@@
Such words are easy to recognize since they have distinct, recurring forms (known affixes, special reduced forms of verbs with incorporated objects), lack internal syntactic markers (e.g. incorporated objects like hal in 1 appear without articles or other modifiers, the verb ‘ire’ meaning “do” is reduced to ‘er’) and are considered single nouns or verbs in terms of parts of speech, as well as single dictionary entries in terms of lemmatization. Due to these multiple levels of complexity, bound groups must first be analyzed and segmented before we can digitize Coptic texts in a way that allows for searchability. Each token can then be lemmatized and tagged to enable structured querying and lexical lookup, or linking to resources such as the Coptic Dictionary Online (CDO, Feder et al. 2018).
@@ 2 Previous work
The most important previous resource we model our work on is the existing UD treebank for Sahidic Coptic Zeldes and Abrams (2018), which contains around 60K words from a range of works in a number of genres. In particular, the Sahidic treebank contains some Biblical material which is in part also available in Bohairic (see Section 3 below). By selecting the same Biblical books and chapters where possible, we are able to conduct some direct comparisons between the dialects which only target parallel passages (see Section 4). At the same time, for texts in the UD Sahidic treebank that are not available in Bohairic, we attempt to select substitutes from similar genres, offering a similar range of language usage.
@@ 3 Data
The text of each of these works is derived from previous digital or print editions. The Life of Shenoute comes from a digital edition created by Hany Takla of the St. Shenouda the Archimandrite Coptic Society, based primarily on the edition published by Leipoldt (1906). Dr. Lydia Bremer-McCollum of the Coptic Scriptorium project [^2] https://copticscriptorium.org/ extracted digitized text using optical character recognition (OCR) from the public domain edition of the Life of Isaac and the public domain edition of the Bohairic Lausiac History, both edited by Amélineau (1887, 1890), followed by manual correction of OCR errors. The Gospel of Mark and 1 Corinthians digital texts come from the Marcion Project; [^3] https://marcion.sourceforge.net/ both ultimately derive from the print edition of Horner (1905). The text of Habakkuk is from the working digital edition in-progress created by the Göttingen Coptic Old Testament project. [^4] https://coptot.manuscriptroom.com/ All of the digital versions were processed by the Coptic Scriptorium project’s natural language processing tools, which have normalized the text, including normalizing variant spellings and expanding any abbreviations; for this paper, all NLP processing has been manually checked by one or more of the authors before treebanking.
@@ 3.1 Inter-annotator agreement
The results show very high agreement, substantially in excess of initial (pre-adjudication) scores in the 80s for the original version of the Sahidic treebank (Zeldes and Abrams, 2018, 199). This is likely due to the facts that the annotators in this case were post-graduate researchers with substantial Coptic annotation experience, as opposed to the novice student scores reported in the Sahidic paper.
@@ 4 Comparing UD Bohairic and Sahidic
To find some of the clearest differences that our data reveals, we extract relative proportions of the dependency relations attested in the Bohairic and Sahidic Coptic treebanks, an excerpt of which is presented in Table 3, which is sorted by the ratio of frequencies.
@@ 4.1 Subject dislocation
One question we can immediately explore using our data and the existing UD Sahidic treebank is whether the constructions are used comparably often in the two dialects. Since segmentation, tagging and parsing guidelines match across the treebanks, we can be confident that all relevant cases can be found using equivalent searches - the results are shown in Table 4.
@@
The table shows that right subject dislocation is much more common in Bohairic, with more hits in total in the smaller Bohairic treebank, and more than double the relative frequency. Looking at the Sahidic data in more detail, it becomes clear that the construction responsible for the discrepancy is nominal subjects in the canonical position between auxiliaries and verbs, as in 4, which is paralleled in Bohairic by 4 – both examples render 1 Corinthians 2:10.
@@
This construction is much rarer in Bohairic, and indicates that the Bohairic data represents a further step in the grammaticalization of the pronominal subject + auxiliary group, which forces subjects to be realized before or after the verbal complex. This tendency is well known from other languages in the Afro-Asiatic family, such as Hausa, which has similar subject + auxiliary structures, but can only realize a nominal subject outside of the verbal complex, with a ‘duplicate’ pronoun mirroring the subject.
@@ 4.3 Focus and preterit marking
The presence of the morpheme glossed as foc in 4.3 indicates that a constituent is focalized in the sentence, in this case ‘water’ (‘I christen you with WATER’ rather than something else).
@@ 5.1 Cross-corpus experiments
Table 5 shows the results for labeled and unlabeled attachment scores on the respective test sets in each setting. The results initially reveal that, unsurprisingly, training and testing across dialects produces very poor results, with LAS and UAS scores around 73 and 62 respectively – the scores are rather comparable in both directions, despite the availability of almost double the data when training on Sahidic. This indicates that the parser is unable to generalize when surface forms vary, since as we noted above, even auxiliaries and prepositions look quite different across dialects.
@@
train LAS UAS LAS UAS
Bohairic 86.349 89.486 62.205 74.633
Sahidic 62.683 73.178 89.760 92.489
Joint 89.929 92.677 88.449 91.602
Balanced 88.989 91.927 86.858 90.628
Table 5: Labeled (LAS) and Unlabeled (UAS) Attachment Scores in each setting when testing on each dialect.
Single dialect models reveal a gap between the smaller Bohairic data and its larger Sahidic counterpart: while Sahidic obtains a LAS of 89.76, Bohairic lags behind with 86.349 (2.5 point difference), with an even starker difference in UAS (92.489 vs. 89.486, about 3 points). Given that the dialects and texts available in them are rather similar, this suggests that more Bohairic data is likely to have an impact in a single-dialect setting.
@@ 5.2 Error analysis
Like English phrasal verbs, some Coptic verbs combine with postposed adverbs to form a unique meaning - for example in 2, cry + out means ‘cry out, call out’ much like in English. In the example, ‘out’ is coincidentally followed by the preposition ‘in’, for the phrase “cry out in a prophetic voice”. However Coptic also has several frequent fixed combination of adverbs with following prepositions, such as {coptic}ebol {coptic}jen, lit. ‘out + in’ but actually meaning ‘out of’ (similar to the English fixed expression ‘out of’ sometimes spelled as ‘outta’). In these cases, our guidelines annotate the second token as fixed, and the first token takes on the expected case label – this ambiguity of adverbs next to prepositions, in which adverbs may belong together with a verb or be part of a multi-word preposition, causes the relatively frequent label confusion in Figure 2. We note that in the interest of comparability with other languages and following UD guidelines (see Ahrenberg 2024 for discussion), the list of fixed expressions of this kind is kept small and is meant to be exhaustive, covering as of writing 25 unique combinations of lemma pairs.
@@ 6 Conclusion
By its nature, this study is based on specific texts which lead to specific results. Although the attempt has been made to select somewhat diverse texts for the corpus, it is always possible that a different selection would have led to different results; we are therefore hopeful that as new data becomes available, additional studies may revisit some of our findings and either validate or relativize these results.
@@ Acknowledgments
We would like to acknowledge work done by members of the Coptic Scriptorium project, the Marcion project and the Coptic Old Testament project, as well as by Hany Takla of the St. Shenouda the Archimandrite Coptic Society, which was instrumental in making the data we annotated in this paper available in digital formats. We also thank the creators of the resources for Sahidic Coptic cited in this paper for making their work available online. All resources developed in this project will be released under open license in the hope that we can contribute and match their generosity.

### v2

@@ 1.1 Coptic
1.2 Bohairic and Sahidic
Expanding digital corpora to include Bohairic texts presents substantial challenges, as existing tools for even basic preprocessing operations, such as word segmentation and lemmatization, let alone syntactic parsing, require tools trained on dialect-specific data (see Section 5 for some evaluation). Although Sahidic and Bohairic are dialectal manifestations of a single language – sharing a broadly consistent grammatical architecture and much of their lexical inventory, their phonological, orthographic and morphosyntactic systems diverge in ways that make NLP tools trained on Sahidic unsuitable for processing Bohairic texts.
At the most elementary level, the two dialects diverge already in their orthographic systems. Unlike earlier Egyptian – written in hieroglyphic and hieratic scripts over the preceding three millennia – Coptic adopted a modified version of the Greek alphabet, with additional characters of Demotic origin (ultimately derived from hieroglyphs) to represent sounds absent from Greek. Both dialects make use of six such additional letters, including the letter hore ({coptic}h1, Unicode U+03E9 lowercase / U+03E8 uppercase) for the voiceless glottal fricative /h/, but Bohairic distinguishes hore /h/ from khei /x/ ({coptic}j , Unicode U+2CC9 lowercase / U+2CC8 uppercase). This distinction is semantically consequential – while the Sahidic word {coptic}eh1rai (ehrai) can confusingly mean both ‘up’ or ‘down’, and can only be disambiguated in context, in Bohairic we see distinct forms that had merged in Sahidic: {coptic}ejrhi (exrēi) ‘down’ and {coptic}eh1rhi (ehrēi) ‘up’.
As an agglutinative language, Coptic combines multiple morphosyntactic elements into units known as bound groups. Following modern editorial conventions, these groups are defined by the presence of a single stressed lexical item at their core Layton (2011). Material that would normally be tokenized into separate words in annotated corpora often appears conflated into a single space-delimited string in Coptic. Such fusion is common, for example, in noun phrases or prepositional phrases, as in 1.2, or in auxiliaries and clitics attaching to verbs, as in 1.2, both examples in the Bohairic dialect.
@@ 1.2 Bohairic and Sahidic
In this paper we use hyphens to split such space-delimited word forms into tokens, which correspond to units that can receive independent parts of speech, such as nouns and verbs, articles and prepositions, etc. Note that such separators do not exist in source texts and are represented in the treebank directly through the tokenization. For words containing derivational affixes as in 1.2, and compound nouns or verbs 1.2, we additionally provide a segmentation into component parts in an additional annotation called MSeg in the conllu format MISC field, following existing practices in UD treebanks. [^1] See https://universaldependencies.org/misc.html#mseg
@@
Such words are easy to recognize since they have distinct, recurring forms (known affixes, special reduced forms of verbs with incorporated objects), lack internal syntactic markers (e.g. incorporated objects like hal in 1.2 appear without articles or other modifiers, the verb ‘ire’ meaning “do” is reduced to ‘er’) and are considered single nouns or verbs in terms of parts of speech, as well as single dictionary entries in terms of lemmatization. Due to these multiple levels of complexity, bound groups must first be analyzed and segmented before we can digitize Coptic texts in a way that allows for searchability. Each token can then be lemmatized and tagged to enable structured querying and lexical lookup, or linking to resources such as the Coptic Dictionary Online (CDO, Feder et al. 2018).
@@ 2 Previous work
The most important previous resource we model our work on is the existing UD treebank for Sahidic Coptic Zeldes and Abrams (2018), which contains around 60K words from a range of works in a number of genres. In particular, the Sahidic treebank contains some Biblical material which is in part also available in Bohairic (see Section 3 below). By selecting the same Biblical books and chapters where possible, we are able to conduct some direct comparisons between the dialects which only target parallel passages (see Section 4). At the same time, for texts in the Sahidic treebank that are unavailable in Bohairic, we select substitutes from similar genres, offering a similar range of language usage.
@@ 3 Data
The text of each of these works is derived from previous digital or print editions. The Life of Shenoute comes from a digital edition created by Hany Takla of the St. Shenouda the Archimandrite Coptic Society, based primarily on the edition published by Leipoldt (1906). Dr. Lydia Bremer-McCollum of the Coptic Scriptorium project [^2] https://copticscriptorium.org/ extracted digitized text using optical character recognition (OCR) from the public domain edition of the Life of Isaac and the public domain edition of the Bohairic Lausiac History, both edited by Amélineau (1887, 1890), followed by manual correction of OCR errors. The Gospel of Mark and 1 Corinthians digital texts come from the Marcion Project; [^3] https://marcion.sourceforge.net/ both ultimately derive from the print edition of Horner (1905). The text of Habakkuk is from the working digital edition in-progress created by the Göttingen Coptic Old Testament project. [^4] https://coptot.manuscriptroom.com/ All of the digital versions were processed by the Coptic Scriptorium project’s natural language processing tools Zeldes and Schroeder (2016), which have normalized the text, including normalizing variant spellings and expanding any abbreviations; for this paper, all NLP processing has been manually checked by one or more of the authors before treebanking.
@@ 3.1 Inter-annotator agreement
The results show very high agreement, substantially in excess of initial (pre-adjudication) scores in the 80s for the original version of the Sahidic treebank (Zeldes and Abrams, 2018, 199). This is likely due to the fact that annotators in this case were post-graduate researchers with substantial Coptic annotation experience, as opposed to the novice student scores reported in the Sahidic paper.
@@ 4 Comparing UD Bohairic and Sahidic
On the lexical level, we can note that of the approximately 2,800 unique words attested in the Sahidic treebank, and around 2100 unique words in the Bohairic treebank, only about 600 are shared, and even among these, identical forms do not always translate to identical meanings. For example while {coptic}et\0 ‘et-’ can be a form of the relative marker in both Sahidic and Bohairic, it can also represent the precursive form meaning ‘after’ in Bohairic. Among the disjoint, dialect-specific vocabulary, many words have corresponding words in the other dialect which differ only due to pronunciation, but some words are totally unique, such as {coptic}baki ‘baki’, which means ‘town’ in Bohairic, but does not exist in Sahidic.
On the syntactic level, we cannot find differences and commonalities as easily, but thanks to the existence of UD trees in both dialects, we can still leverage annotated data in a straightforward way. To find some of the clearest differences that our data reveals, we extract relative proportions of the dependency relations attested in the Bohairic and Sahidic Coptic treebanks, an excerpt of which is presented in Table 3, which is sorted by the ratio of frequencies.
@@ 4.1 Subject dislocation
One question we can immediately explore using our data and the existing UD Sahidic treebank is whether the constructions are used comparably often in the two dialects. Since segmentation, tagging and parsing guidelines match across the treebanks, we can be confident that all relevant cases can be found using equivalent searches – the results are shown in Table 4.
@@
The table shows that right subject dislocation is much more common in Bohairic, with more hits in total in the smaller Bohairic treebank, and more than double the relative frequency. The bottom of the table shows results for the same queries, restricted to only the parallel data available in both corpora – these numbers are slightly less reliable since they are based on less material, but should be expected to be much closer since they derive from the same chapters of the Bible. Here too, the gap remains very substantial, despite the fact that if anything Biblical style should represent a conservative and less free or colloquial style. [^6] An anonymous reviewer has also asked about comparable breakdowns by genre, but we feel that for some of the texts, comparisons would be hard to establish, and these would also split the already small treebank to the point where counts would be substantially less reliable.
Looking at the Sahidic data in more detail, it becomes clear that the construction responsible for the discrepancy is nominal subjects in the canonical position between auxiliaries and verbs, as in 4, which is paralleled in Bohairic by 4 – both examples render 1 Corinthians 2:10.
@@
This construction is much rarer in Bohairic, and indicates that the Bohairic data represents a further step in the grammaticalization of the pronominal subject + auxiliary group, which forces subjects to be realized before or after the verbal complex. This tendency is well known from other languages in the Afro-Asiatic family, such as Hausa, which has similar subject + auxiliary structures, but can only realize a nominal subject outside of the verbal complex, with a ‘duplicate’ pronoun mirroring the subject – a pronominal TAM marker within the verbal complex is mandatory (see Crysmann 2012, 331, and Hartmann 2006 on fronting in Hausa).
@@ 4.3 Focus and preterit marking
The presence of the morpheme glossed as foc in 4.3 indicates that a constituent is focalized in the sentence, in this case ‘water’ (‘I christened you with WATER’ rather than something else).
@@ 5.1 Cross-corpus experiments
Table 5 shows the results for labeled and unlabeled attachment scores on the respective test sets in each setting, along with train and test partition sizes in thousands of tokens. The results initially reveal that, unsurprisingly, training and testing across dialects (red numbers) produces very poor results, with LAS and UAS scores around 73 and 62 respectively – the scores are rather comparable in both directions, despite the availability of almost double the data when training on Sahidic. This indicates that the parser is unable to generalize when surface forms vary, since as we noted above, even auxiliaries and prepositions look quite different across dialects.
@@
(11K tokens) (10.3K tokens)
train (tokens) LAS UAS LAS UAS
Bohairic (16.5K) 86.349 89.486 62.205 74.633
Sahidic (35.8K) 62.683 73.178 89.760 92.489
Joint (52.3K) 89.929 92.677 88.449 91.602
Balanced (36K) 88.989 91.927 86.858 90.628
Table 5: Labeled (LAS) and Unlabeled (UAS) Attachment Scores in each setting when testing on each dialect. Within and across dialect scores are in green and red respectively. The best settings for each dialect are bolded.
Single dialect models (green numbers) reveal a gap between the smaller Bohairic data and its larger Sahidic counterpart: while Sahidic obtains a LAS of 89.76, Bohairic lags behind with 86.349 (2.5 point difference), with an even starker difference in UAS (92.489 vs. 89.486, about 3 points). Given that the dialects and texts available in them are rather similar, this suggests that more Bohairic data is likely to have an impact in a single-dialect setting.
@@ 5.2 Error analysis
Like English phrasal verbs, some Coptic verbs combine with postposed adverbs to form a unique meaning – for example in 2, cry + out means ‘cry out, call out’ much like in English. In the example, ‘out’ is coincidentally followed by the preposition ‘in’, for the phrase “cry out in a prophetic voice”. However Coptic also has several frequent fixed combination of adverbs with following prepositions, such as {coptic}ebol {coptic}jen, lit. ‘out + in’ but actually meaning ‘out of’ (similar to the English fixed expression ‘out of’ sometimes spelled as ‘outta’). In these cases, our guidelines annotate the second token as fixed, and the first token takes on the expected case label – this ambiguity of adverbs next to prepositions, in which adverbs may belong together with a verb or be part of a multi-word preposition, causes the relatively frequent label confusion in Figure 2. We note that in the interest of comparability with other languages and following UD guidelines (see Ahrenberg 2024 for discussion), the list of fixed expressions of this kind is kept small and is meant to be exhaustive, covering as of writing 25 unique combinations of lemma pairs.
@@ 6 Conclusion
By its nature, this study is based on specific texts which lead to specific results. Although the attempt has been made to select somewhat diverse texts for the corpus, it is always possible that a different selection would have led to different results. In particular, the inclusion of translated texts, such as material from the Bible, is not ideal for some types of research, but as is often the case in resources for historical languages with limited attestation, this is somewhat inevitable. We are hopeful that as new data becomes available, additional studies may revisit some of our findings and either validate or relativize these results.
@@ Acknowledgments
This work was made possible by a grant from the National Endowment for the Humanities Preservation and Access Humanities Collections Reference and Resources Program (PW-290519-23). We thank our postdoctoral fellow, Lydia Bremer-McCollum, for digitizing some of the text used in this paper via OCR, and Wolfgang Jentner and the Data Institute for Societal Challenges at the University of Oklahoma for their server support. We would also like to acknowledge work done by other members of the Coptic Scriptorium project, the Marcion project and the Coptic Old Testament project, as well as by Hany Takla of the St. Shenouda the Archimandrite Coptic Society, which was instrumental in making the data we annotated in this paper available in digital formats. All resources developed in this project will be released under an open license in the hope that we can contribute and match their generosity.